OpenStudy (anonymous):

please check if differential equations are solved correctly dy/dx – 4y = e˟ given that y(0) = 0 using dy/dx + Py = Q P = -4 & Q = e˟ Integrating factor = e˄∫-4dx e¯⁴˟ solution: (e¯⁴˟)y = ∫ (e¯⁴˟ )( e˟) dx (e¯⁴˟)y = ∫ (e¯³˟ ) dx (e¯⁴˟)y = -¼ e¯⁴˟ + C y = -¼eº +C(e⁴˟) y = -¼ +C(e⁴˟) y(0) = -¼ +C(e⁴˟) 0 + ¼ = C(eº) C = ¼ y = -¼ + ¼ e⁴˟

OpenStudy (phi):

I would first solve the homogeneous equation y' -4y = 0 using the characteristic equation m-4=0 then use the "method of undetermined coefficients" to find the particular solution in this approach you assume y = A e^x plug that into the diff eq and solve for A

OpenStudy (phi):

if we use the integration factor, you are ok up through solution: (e¯⁴˟)y = ∫ (e¯⁴˟ )( e˟) dx (e¯⁴˟)y = ∫ (e¯³˟ ) dx the integral of (e¯³˟ ) dx is -1/3 e¯³˟ NOT -¼ e¯⁴˟ apparently you are thinking of the power rule of integration. but the integral of e^x dx is e^x in this case you want to integrate \( -\frac{1}{3}e^{-3x} (-3) dx \)

OpenStudy (anonymous):

Thanks, fixed it as follows: Solution: (e¯⁴˟)y = ∫ (e¯⁴˟ )( e˟) dx (e¯⁴˟)y = ∫ (e¯³˟ ) dx (e¯⁴˟)y = -⅓ e¯³˟ + C y = -⅓ (e⁴˟)(e¯³˟) + e⁴˟ C y = -⅓ (e˟) +C(e⁴˟) y(0) = -⅓ (e˟) +C(e⁴˟) 0 + ⅓ = C(eº) C = ⅓ y = -⅓ + ⅓(e⁴˟)

OpenStudy (phi):

yes, you can always check it by putting it into the original differential equation

OpenStudy (phi):

although your last line left out e^x from the first term

OpenStudy (anonymous):

thanks again.